Calculate the pH buffer which will permit 0.1 all but 0.1mg of Zn to be precipitated from ZnSO4 solution when saturated with H2S. how many grams of formic acid and sodium formate should be added to the solution to prepare this buffer?
1
Expert's answer
2020-06-09T05:38:53-0400

pH = pKa + log (Cbase/Cacid)

pH = 1.8x10-4 + log (Cbase/Cacid)

pH = pKa + log(m(HCOONa)/68 / m(HCOOH)/46)

m(HCOONa)/m(HCOOH) * 0.676 = 10pH-pKa

If any of two quantities m (HCOONa) or m (HCOOH) is provided, the one may calculate how many grams of formic acid and sodium formate should be added to the solution to prepare the buffer.